Solve the system. (Please show all work.) I will be rewriting it in operator notation as shown below

\frac{d^2x}{dt^2}-4x-2\frac{dy}{dt}+y=t \Rightarrow (D^2-4)x+(1-2D)y=t

2\frac{dx}{dt}+x+\frac{d^2y}{dt^2}=0 \Rightarrow (2D+1)x+D^2y=0
Answer #1

Ans Given that Equation can be re-written as (0²-4) x + (1-2D) y = t where Doa Equation can be re-written as (1 +20)x + 2y =(m²-1) (m2+1) = 0 => m2-1-o , m?+1 to > m=# , mrti Complementary function will be C.F. = Get + Get & Cz Cost + 4 sint where GMow particular integral (P.I.) = 1 A+ 2) (1-DA) - P.J. = (1-DA) (t + 2) = (1 + Dt + Do + ...) (t +2) = (t + 2) +0 = A +2 P.I.